OpenStudy (anonymous):

A tourist group with 5 people uses the water that they took over 14 days. Two more join in. How many days would it take for the water to finish ?

OpenStudy (anonymous):

How long would the water last now?

OpenStudy (anonymous):

Use this equation: 14*5/7 =70/7 =10. It would take a tourist group of 7 people 10 days to finish the water.

OpenStudy (anonymous):

Unless - this is a trick question and the water's already gone when the other 2 join. :)

OpenStudy (anonymous):

nope you got it! Thanks a lott :)

OpenStudy (anonymous):

no problem. :)